### v1

# A Mathematical Framework for Custom Reward Functions in Job Application Evaluation using Reinforcement Learning

Abstract
Conventional Applicant Tracking Systems (ATS) tend to be inflexible keyword-matchers, and deny gifted candidates a role due to a few minor semantic mismatches. This article describes a new two-step process to design a more refined resume evaluation model based on a small language model (¡600M parameters) that is finetuned using GRPO on a custom reward function. To begin with, Supervised Fine-Tuning (SFT) was used to build a solid baseline model. Second, this SFT model was also optimized with the help of Reinforcement Learning (RL) through GRPO under the guidance of a new, multi-component reward function that can holistically assess candidates beyond simple keyword matching. We indicate that the RL application presents a critical problem of reward hacking due to the initial experiments of aggressive penalties, which produces faulty, excessively negative model behaviors. We have overcome this challenge by refining the reward function repeatedly and training hyperparameters into a stable “gentle polishing process” of the reward function. Our resulting GRPO-polished model demonstrates significant real-world efficacy, achieving a final accuracy of 91% on unseen test data. The model shows a strong ability to correctly identify qualified candidates (recall of 0.85 for the ‘SELECTED’ class) while also showing exceptional precision (1.0), confirming its reliability. These results indicate that a properly executed, two-step fine-tuning procedure can indeed effectively refine a small language model to be able to conduct fine-tuned and human-like candidate scoring, overcoming the drawbacks of both traditional ATS and naive RL usage.
keywords:
Applicant Tracking System , Fine-tuning , Small Language Model , Reinforcement Learning , Custom Reward Function
footnotetext: Work done during an internship at O6AI LABS.
Emails: {shreyansh.jain, madhav, shreya, pranav, dishaa.l, naren, akash.anandhan}@o6ai.com
footnotetext: * Core Contributors
1 Introduction
The automation of the industries in the era of Artificial Intelligence makes the field of recruitment one of the most significant areas of development. The screening process should have been automated long ago due to the fact that there are thousands of applicants per position, thus making manual screening of applicants obsolete, and forcing the use of Applicant Tracking Systems (ATS). ATS tools are however susceptible to the following weaknesses: they are somewhat reliant on the keyword filtering and fail to consider such critical parameters as contextual skills, work-related experience, and the quality of education. The end result is loss of highly qualified candidates and unfair dismissal which cause opportunity costs on part of organizations.
Our research proposes an intelligent evaluation pathway powered by Small Language Models (SLMs) to overpower them. In comparison to enormous Large Language Models (LLMs), SLMs—usually in the 200 to 600M range—are cheaper, predictable and better suited to the purpose of recruitment activities, which they are able to perform in a fine-grained fashion. Smaller models reduce hallucinations and adapt better to limited data, making them well-suited for agentic systems [4].
Our study has made a contribution: the two phase training regime. The Supervised Fine-Tuning (SFT) is the first, where the SLM is trained on the grammar of recruitment-related job ads, resumes and key skills indicators to learn to decipher job ads. Then, a custom reward is applied to reinforcement learning to optimize by Group-Relative Policy Optimization (GRPO) [7] to take into account a wide range of skills and experience education levels to align with the recruiter judgement. This allows comprehensive candidate analysis to go beyond traditional ATS.
The contributions of the paper include: a novel resume evaluation pipeline that does not use a strict keyword matching threshold, but does rely upon ranking the set based on AI based relevance scoring; a demonstration of the effectiveness of SLM in a highly specific domain through domain specific fine-tuning; and being the first reported use of GRPO in HR technology, in matching the AI output to human opinion. Trained on a dataset of nearly 3,000 resumes across 337 RL steps, our results demonstrate the potential for this approach to serve as a powerful tool against inequitable and inefficient talent acquisition.
2 Literature Survey
Traditionally, the recruitment technology was a compromise between efficiency and effectiveness of automated screening and human evaluation respectively. Traditional Applicant Tracking System (ATS) were the most popular and were not immune to criticism due to their reliance on strict keyword-matching. Van Esch et al. [1] note that this weakness can cause them to unwillingly filter out qualified applicants because of tiny semantic signals which can also be observed in recent research problems on the topic of algorithmic bias in hiring by Albaroudi et al. [2].
This is the constraint that has seen the pursuit of more advanced ways of machine learning capable of extending the semantic associations. Previous works focused on enabling a more linguistic understanding of resumes and job descriptions using Transformer-based approaches and deep learning embeddings like BERT and GPT [3]. Despite the fact that these neural embedding methods have significantly improved the matching performance, there is a tendency toward the expansion of models, which has questioned the cost of calculation.
More recently, such a change has been in the spotlight with Belcak et al. [4] reporting that such a shift has occurred, and the focus has shifted to the potential of Small Language Models (SLMs) to possess an appealing balance between performance and efficiency of domain-specific tasks like resume evaluation. This challenges the old-time belief that bigger models are always better and that domain-specific fine-tuning is important.
Simultaneously, Reinforcement Learning from Human Feedback (RLHF), firstly introduced in the literature by Stiennon et al. [6] and Ouyang et al. [5], has become an effective framework of creative language model-dependent model tuning. Recent methods, like Group Relative Policy Optimization (GRPO), proposed by Shao et al. [7], have demonstrated their potential of enhancing the cost-effectiveness of this alignment process.
A major difficulty in the application of RL however is the phenomenon of reward hacking, in which a model manipulates the reward term’s weaknesses without achieving the desired objective. This may cause unwanted habits, such as the acquisition of an over-pessimistic bias [8, 9], as explained in recent articles. This concern indicates the basic constraints and unfinished issues in the sphere, according to Casper et al. [10].
Notwithstanding all these developments, there are a few gaps in research. Usage of contemporary RL methods such as GRPO with SLMs in human resource sphere has been little explored. Moreover, there is a lack of practical research of the issue of reward hacking in this particular environment. Current systems do not tend to deliver a holistic evaluation of the candidate and the issue of matching model performance and the multi-objective, human recruiter remains. These gaps are filled out by our work, which initially provides the first reported application of GRPO to fine-tune an SLM to resume evaluation, a concrete case study of what we believe are the challenges of reward hacking, and a refined, multi-component reward function as a practical solution.
3 Proposed Methodology
The proposed AI resume evaluation agent is trained in two phases. Supervised Fine-Tuning (SFT) is first applied to get a baseline idea of the task, and then Generative Reward Policy Optimization (GRPO) is applied to refine the reasoning of the model to match expert-heuristic reasoning. The data was artificially created so that there was an equal representation of approval and rejection classes. Resume and job description templates and logical rules were used to create candidate resumes and job descriptions programmatically in order to simulate realistic recruitment conditions, but eliminate privacy concerns in real resumes. The dataset, whilst artificial, was to be internal consistent (skills, experience and outcomes) to offer a testbed of a valid evaluation.
3.1 Model Selection and Configuration
In the case of the base model, we picked unsloth/Qwen2-0.5B-Instruct-bnb-4bit as it is efficient and the best performance on the baseline. We used 4-bit quantization using the Unsloth library and PEFT through LoRA (rank = 16, $\alpha$ = 32), which allows us to perform efficient adaptation without refining all the parameters. The stage of SFT (3,000 samples (90% train, 10% validation)) was provided in the format of prompts where the model is being asked to perform as an HR expert and provide a response in the form of a JSON object with a score and binary status (SELECTED or REJECTED).
There were two epochs of training at a learning rate of $2\times 10^{-4}$ (linear scheduler) and adamw-8bit optimizer. The optimal batch size was 8 and the per-device batch size was 2 and the number of gradient accumulation was 4. This configuration offered effective, memory-conscious training and also guaranteed consistent gradient updates as well as avoiding overfitting.
3.2 GRPO Refinement Phase
The second step would improve the SFT-tuned model, which would involve improving the quality and logical consistency of the evaluations. This is done by optimizing the policy of the model over a hand-crafted, multi-component reward function that is intended to promote more human-like, fined-grained reasoning and positively discourage the act of reward hacking.
3.2.1 Reward Formulation
The center of the GRPO stage is a reward function as in Eq. 1 which gives a single holistic feedback signal in a weighted combination of four criteria. This interdisciplinary nature is the main tool to combat reward hacking since the model needs to meet many, even conflicting, goals in order to reach a high reward and cannot be able to rely on a single, easy measure.
$\text{Reward}=\sum_{i=0}^{N}(W_{i}*S_{i})$ (1)
Where: (1) $N$ = number of evaluation criteria (here, $N=4$); (2) $W_{i}\in[0,1]$ weight assigned to criterion, subject to $\sum_{i=0}^{N}W_{i}=1$; (3) $S_{i}$ = score assigned to criterion determined by task-specific rules; (4) $i$ is the index of the criteria.
The reward formulation proposed is based on the principle of weighted linear combination, which is similar to artificial neural network feature activation aggregation by weighted summations. The evaluation criteria have a proportional contribution to the total reward, so that the contribution of each factor is not dominant without the weight being explicitly specified.
Final Reward Calculation.
The weights and score ranges shown in Table 1 were determined through iterative empirical tuning to maximize model stability and alignment during GRPO training. Multiple configurations were evaluated, and the final values were selected based on the best balance of reward sensitivity, classification performance, and avoidance of reward hacking.
Table 1: Parameter Configuration
Parameter Base Score Range Scoring Rules
Weight ($W_{i}$)
Status Correctness 0.40 $S_{i}\in[-2,2]$ +2: TP; 0: TN; -1: FP; -2: FN
Score Accuracy 0.20 $S_{i}\in[-1,1]$ +1: Matches expected score; 0:
Consistent with score; -1:
Invalid/illogical
Skills Matching 0.20 $S_{i}\in[-1,1]$ +1: $\geq$75% skill match; 0: 40–
74% match; -1: ¡40% or no data
Experience 0.20 $S_{i}\in[-1,1]$ +1: Score & status match level;
Evaluation 0: Partial alignment; -1:
Misaligned
Early experiments showed that overly aggressive penalty ranges caused pessimistic model behavior, whereas more moderate configurations led to stable and human-aligned policy updates. The chosen formulation reflects the most stable configuration observed during tuning.
$R=\sum_{i=1}^{4}(W_{i}*S_{i})$ (2)
3.3 Training Setup Notes
GRPO training setup is carefully planned not just to be policy-optimal, but also to provide a solid defense against reward hacking, the behavior where a model uses the reward function to take advantage of the policy to score highly with nonsensical or undesirable outputs. The same 3,000 samples are used in this phase as in the SFT phase. This is a standard and intentional procedure, the idea of GRPO is not to learn anything new based on the labels of the dataset, but to optimize the reasoning policy of the model.
The multi-faceted reward function itself is our primary preventative tool, however, the training dynamics are the second level of defense which is critical. It starts with loading a SFT adapter, restoring the policy with a task-conscious, stable baseline. The set-up is then adjusted to a softer polishing instead of hard optimization. Very small learning rate of $2\times 10^{-6}$ is used to make small and consistent policy changes.
It is important to note, and in contrast to the SFT phase, the GRPO training loop does not use an evaluation dataset deliberately. The refinement in the policy is solely informed by the reward signal produced by the training samples since the conventional measures of validation, such as accuracy, are not very useful in this optimization scenario. The model is trained only one epoch, or 337 training steps on our data. This short time limit is an intended option to restrict its exposure to the reward landscape and decrease the risk of over-optimizing.
The simplest system to directly overcome this policy drift is the application of KL-divergence regularization, whose regulation is determined by the beta parameter, which should be 0.1. This regularization punishes the model when the output policy of the model becomes too different to the original SFT policy. This functionally restrains the model to a space of realistic and sensible solutions that it is trained on during SFT and discourages it to produce bizarre and high-reward outputs.
4 Results
The experimental analysis shows that the model performance is dramatically improved throughout the two-phase training process as both the Supervised Fine-Tuning (SFT) and Generative Reward Policy Optimization (GRPO) stages provide positive results.
The SFT phase, which was done in two epochs on 2,700 training samples in total, managed to achieve a powerful baseline model. This training used a useful batch size of 8 and fine-tuned 8,798,208 LoRA parameters, 1.75% of the total 502,830,976 parameters of the model. The training and validation losses to each other converged successfully as shown in Figure 1. The training loss reduced very fast initially, and the final value was around 0.28, which means that model was able to learn basic structure and format of resume evaluation task. Validation loss tended to follow training loss, as it reached a similar value, which proves that the model was not overfitting and was able to predict on the data that it had not seen before.
Figure 1: Training and Validation Loss Curves for SFT
A GRPO optimization, which was performed over one epoch and took 337 steps, had more interesting and significant dynamics. The training loss reduced drastically by 97.3 percent (the initial training loss was 4.9380 and now it is 0.1337). This drastic decrease is an indication of the utility of the custom reward feature in the optimization of the policy of the model. This is also supported by the reward metric that evened out to a final outlook of 0.0330-0.0330 meaning that the assessments of the model became more corresponding to our preferred requirements. At the same time, the KL divergence came to a final value of 0.34767, which demonstrates that the policy was optimized and stabilized without losing the knowledge acquired in the course of the SFT stage. Table 2 provides a summary of these important metrics.
Table 2: Comparing Pre-trained and Post-trained Model
Method Initial Final Loss Final KL
Loss Loss Reduction Reward Divergence
SFT 0.4070 0.2796 31.3%
GRPO 4.9380 0.1337 97.3% 0.0330-0.0330 0.34767
KL divergence metric is used to evaluate the extent of divergence of the policy of the model as compared to that of the original SFT policy. The trend has been equivalent to the training loss with an initial steep decline and then leveled to the low value as shown in Figure 2. Such is the optimal behavior: it demonstrates that the model is making serious, constructive changes to its policy very early (large initial KL) but soon adopts a sophisticated state, without wandering too far out of the original knowledge it gained in the process of SFT (small final KL of 0.34767). This proves the fact that the training was balanced, and the so-called leash that the KL penalty offers worked in averting the collapse of the policy.
Figure 2: GRPO Training Loss over 337 steps
Subsequent examination of the training logs will show the consistency of the generation process. During the GRPO phase, the average length of the responses generated was always maintained between 60 and 90 tokens, and the average length of terminated responses was 30 to 50 tokens. This shows that the model had been trained to generate outputs of the appropriate and constant length. Moreover, the clipped ratio, indicating the share of policy changes, was changing, but tended to be in the 0.1 to 0.2 interval, indicating consistent and restrained policy changes during the course of training.
4.1 Comparative Performance on Test Data
To quantify the actual effects of the GRPO phase, both SFT-only model and the final GRPO-refined model were tested on the held-out test set of 104 unseen samples. GRPO-refined model performed better in all the key metrics compared to the SFT-only baseline. According to Table 3, the GRPO model had a better classification accuracy (91.4% vs. 89.4%), and better F1-Score on the important class, i.e. selected (0.92 vs. 0.90). This validates that the phase of policy alignment did not only enhance the internal logic within the model but also carried over to the more precise final decisions.
Table 3: Comparative Performance on the Test Set
Metric SFT-Only GRPO-Refined Improvement
Model Model
Overall Accuracy 89.4% 91.4% +2.0%
F1-Score 0.9043 0.9204 +1.8%
Mean Absolute Error (MAE) 16.05 15.47 $3.6\%$
RMSE 19.81 19.49 $1.6\%$
Figure 3 provides a visual comparison of the performance metrics between the SFT-only and GRPO-refined models.
Figure 3: Bar chart comparing SFT-only and GRPO-refined model performance
When analyzed in granular detail through the confusion matrices shown in Figure 4, there was a crucial improvement of the decision-making of the GRPO model. The SFT model wrongly identified 2 REJECTED candidates as SELECTED. GRPO model removed these false positives altogether and the number of false positives dropped to zero, at the same time, the number of successfully identified ‘REJECTED’ candidates (True Negatives) increased to 43.
Figure 4: Confusion matrices for SFT and GRPO models
This outcome is a direct success of the reward function’s design. By penalizing incorrect classifications, the GRPO phase created a more discerning model that is less likely to pass unqualified candidates to the next stage, thereby improving the efficiency of the hiring pipeline.
5 Conclusion
This study manages to prove that a two-stage SFT and GRPO pipeline can be used to convert a small language model into an advanced resume judging system, despite traditional ATS being inflexible. Our last model gave 91 percent accuracy on unknown data and its ability to predict the chosen candidates was 98 percent, confirming its usefulness in the real world. The total success rate increase of the GRPO phase was 2.0 per cent, but the implications of these gains in practice are enormous. As an example, the refined model has zero false positive in the test set, the number of which was 2. In a practical hiring pipeline, this is a major time and cost savings as it means that unqualified applicants do not go through more resource-intensive steps.
Our main contribution is our multi-component reward function, which makes the model consistent with the complex business logic, most importantly, false negatives, and our training strategy of gentle polishing proved to be useful in reducing reward hacking. This paper introduces a computationally-efficient model building system to create expert models with nuanced and human-like reasoning, without the need to scale to massive architectures. The model can be further improved in the future to address its practical use further by allowing the model to identify candidacies of ambiguous grey areas and mark them as subject to manual review, to become a collaborative worker of decision-support.